OpenStudy (anonymous):

Solve the problems: 1. 2/3=1.2/x a)x=18 b)x=1.8 c)x=0.2 d)x=10 2. 4/5=8/x+2 a)x=4 b)x=.5 c)x=2 d)x=8

OpenStudy (anonymous):

@Amy143

OpenStudy (anonymous):

Oh, sorry honey, I have no clue how to do this one.

OpenStudy (anonymous):

thx :p

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

these are proportions... 2/3 = 1.2/x cross multiply (2)(x) = (1.2)(3) 2x = 3.6 x = 3.6/2 x = 1.8 any questions about this one ?

OpenStudy (anonymous):

all u have to do is cross multiply?

OpenStudy (anonymous):

yes...and solve for the missing variable

OpenStudy (anonymous):

oh ok but i'm still confused ლ(́◉◞౪◟◉‵ლ)

OpenStudy (anonymous):

what exactly are you confused about ?

OpenStudy (anonymous):

well math is hard for me so I still don't know how to find the missing variable even though u just said cross multiply.. I'm not very good at math :(

OpenStudy (anonymous):

shall we work on the second problem ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I have 1 question...is the problem 4/5 = (8/x) + 2 or is it 4/5 = 8/(x + 2) ??

OpenStudy (anonymous):

well it's 8 over x+ 2

OpenStudy (anonymous):



OpenStudy (anonymous):

4/5 = 8/x + 2 cross multiply 4(x + 2) = (5)(8) 4x + 8 = 40 4x = 40 - 8 4x = 32 x = 8

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

No I'm starting to get it now. Thank you very much!

OpenStudy (anonymous):

no problem...glad to help :)

OpenStudy (anonymous):

:)